Write the quadratic function in vertex form.

y = x2 + 16x + 74

y = (x + 8)2 + 10

y = (x + 8)2 - 10

y = (x - 8)2 + 10

y = (x - 8)2 - 10


Help me please
1
Expert's answer
2014-06-09T03:17:42-0400
y = x2 + 16x + 74 = x2 + 16x + 64 + 10
Using the formula (a+b)2 = a2 + 2ab + b2 we get
y = x2 + 16x + 64 + 10 = (x + 8)2 + 10.